Jess and david took an 800-mile driving trip.jess drove 200 miles more than david did on the trip.how many miles did david drive?
Answer by addingup(3677)

j+d= 800
j= d+200 Substitute for j in the equation above:
d+200+d= 800
2d+200=800
2d= 600
d= 300 This is how many miles David drove
Proof:
David drove ...................300
Jess drove 200 more: 300+200= 500
_ _ _ _ _ _ _ _ _ _ _ _ _ _ _ --------
Total..........................800 We have the right answer